OpenStudy (anonymous):

a student earned on tests the scores of 88, 75, 89, and 81. what is the lowest score the student must get on the next test to keep an average of 80?

OpenStudy (anonymous):

can you help on this @jim_thompson5910

jimthompson5910 (jim_thompson5910):

(88+75+89+81+x)/5 = 88 solve for x

OpenStudy (anonymous):

is that suppose to be divided by 5 ? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

yes, you add up all 5 scores, then you divide by 5 to average them

jimthompson5910 (jim_thompson5910):

sorry, the equation should be (88+75+89+81+x)/5 = 80

jimthompson5910 (jim_thompson5910):

misread but I fixed it

OpenStudy (anonymous):

so i got 333.. so then i divide by 5 and i got 66.6 @jim_thompson5910

jimthompson5910 (jim_thompson5910):

sorry I changed the equation to (88+75+89+81+x)/5 = 80 (misread initially)

jimthompson5910 (jim_thompson5910):

(88+75+89+81+x)/5 = 80 88+75+89+81+x = 80*5 333+x = 400 x = ???

OpenStudy (anonymous):

67 ? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

yep

OpenStudy (anonymous):

ok thanks ... @jim_thompson5910

jimthompson5910 (jim_thompson5910):

np